OpenStudy (camerondoherty):

\[f(x) = x^3 + 4x^2 - x - 4\] What are the steps you would follow to graph f(x)? Describe the end behavior of the graph of f(x).

OpenStudy (camerondoherty):

@surjithayer Do you know anything about this? Im stumped. :/

OpenStudy (anonymous):

\[f \left( x \right)=x^3+4x^2-x-4=x^2\left( x+4 \right)-1\left( x+4 \right)=\left( x+4 \right)\left( x^2-1 \right)\] \[=\left( x+4 \right)\left( x+1 \right)\left( x-1 \right)\] when x=0,f(x)=-4 f(x)=0, when x=-4,-1,1 \[f \prime \left( x \right)=3x^2+8x-1,\] put f'(x)=0,find x and then f(x) at these points at these points tangent is parallel to x axis f(x) is increasing if f'(x)>0 and find the interval for x. f(x) is decreasing if f'(x)<0 and find the interval for x. keeping these points in mind you can draw the graph considering f(x)=y

OpenStudy (camerondoherty):

Just to clear something up So using this equation \[f \prime \left( x \right)=3x^2+8x-1,\]You sub the values for x, and find f(x)? Then you draw the graph accordingly? c:

OpenStudy (camerondoherty):

Sorry that took so long to type I had it there, but the page kept going off somehow and I wasn't able to post it

OpenStudy (anonymous):

some times it is difficult to understand if it is increasing or decreasing

OpenStudy (camerondoherty):

Wait, but you do find f(x) and if it is greater than 0 it is increasing?

OpenStudy (anonymous):

no, I have written \[f \prime \left( x \right)>0\] if it is increasing.

OpenStudy (camerondoherty):

Oh I see, Thank You so much c:

OpenStudy (anonymous):

yw